OpenStudy (anonymous):

Solve the equation sin(2t) + cost= ¼ on the interval [0, 2π]. Express answer(s) to three decimal places.

OpenStudy (anonymous):

I graphed y=sin(2t)+cost and then y=1/4, but they are coming out as straight lines. did I do something wrong?

OpenStudy (anonymous):

expand sin(2t) using trig identities. Do you have the identities infront of you now? I dont know them off by heart

OpenStudy (anonymous):

No

OpenStudy (anonymous):

2sintcost +cost = 1/4

OpenStudy (anonymous):

@hartnn any idea how one would solve this question?